OpenStudy (anonymous):

Christopher has scored at least 1 goal in 12 of his 15 soccer games this season. What is the experimental probability that he will score in the next game?

OpenStudy (shadowreaper810):

Total Outcomes observed are 15. The Outcomes that were favourable(scoring a goal) are 12 so Probability=12/15 P(E)=12/15=4/5=0.8